OpenStudy (curry):

sin 3x= 3cos^2(x)sinx - sin^3(x) prove

OpenStudy (amistre64):

well, first id wanna split that 3x into parts

OpenStudy (amistre64):

sin(2x+x) then split the results again

OpenStudy (amistre64):

sin(2x+x) = sin(2x)cos(x)+sin(x)cos(2x) = sin(x+x)cos(x)+sin(x)cos(x+x) etc

OpenStudy (amistre64):

with any luck it simplifies down to that

OpenStudy (curry):

it doesnt simplify down to that

OpenStudy (amistre64):

what did you end up with?

OpenStudy (curry):

4sincos^2-sin

OpenStudy (amistre64):

sin(x+x)cos(x)+sin(x)cos(x+x) 2sin(x)cos^2(x) + sin(x)cos^2(x)-sin^3(x) 3sin(x)cos^2(x) -sin^3(x)

OpenStudy (amistre64):

you sure?

OpenStudy (curry):

um

OpenStudy (curry):

nvm u got it right i used ther wrong identity